Deduce gas laws from kinetic gas equation with proofs

Grade:11

1 Answers

Vikas TU
7 years ago
Kinetic gas eqn.
PV = nRT
and K = 3KT/2.
  1. Boyle’s law => PV =constant that is at constant temperature,    
  2. P1V1 = P2V2
  3. Charle’s Law => Presuure is constant
  4. that is = V1/T1 =V2/T2